OpenStudy (anonymous):

what is its factorization 2x^2+8x+8=2(x+2)^2 is that correct ?

OpenStudy (anonymous):

yes, it is correct.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

\[2x ^{2}+8x+8=2(x ^{2}+4x+4)=2(x+2)(x+2)=2(x+2)^{2}\] So it is correct.